a gas is enclosed in a vessel of volume V at a pressure P.it is being pumped out of the vessel by means of a piston pump witha stroke volume V/100 . What is the final pressure in the vessel after n strokes of the pump ?


Assume no change in temperature during pumping out of gas .

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Ankit

PV=NRT

N is the number of molecile of gas initially

given T=constant

V is also constant

so

P/N =constant

after one stroke molecules of gas remain in cylender is N(1-1/100)

after 2 stroke molecules of gas remain in cylender is N(1-1/100)2

after n stroke molecules of gas remain in cylender is N(1-1/100)n

so P/N =P'/N(1-1/100)n

or P'=P(1-1/100)n


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin